OpenStudy (anonymous):

A Rocket is shot horizontally off a cliff that is 49 m high. If its initial velocity is 125 m/s. How long will the rocket be in the air? How far away from the cliff will the rocket land on the ground?

OpenStudy (anonymous):

y = 1/2gt^2 =- 1/2(-9.8)t^2 = /t = sqrt{2y/g} = 2(49)/9.8 = 98/9.8 = sqrt(10) = A) 3.2seconds x= Vxo * t= (125ms)(3.2seconds) = 400m